OpenStudy (anonymous):

how can find of value sin18

OpenStudy (agent0smith):

Do you mean without a calculator...?

OpenStudy (anonymous):

no i mean that total method of solving

OpenStudy (agent0smith):

I'm not sure what you mean by "total method of solving"... care to explain?

OpenStudy (anonymous):

to aap jante kya hai

OpenStudy (agent0smith):

Google translate didn't help me... except point out that you might be speaking Finnish.

OpenStudy (anonymous):

lol..

OpenStudy (anonymous):

He said what do you mean...

OpenStudy (agent0smith):

So he said what do you mean after I asked him what he meant? so we were basically stuck in a loop of not understanding each other.

OpenStudy (anonymous):

I think its a language barrier...

OpenStudy (agent0smith):

You don't say? Here i thought "to aap jante kya hai" was perfect english :P

OpenStudy (anonymous):

I know sounds so pure!

OpenStudy (agent0smith):

haha :)

OpenStudy (anonymous):

Think hes from India, they dont let you use calculators there so he was probably like huh????

OpenStudy (agent0smith):

haha oh i see

OpenStudy (amistre64):

\[sin(x)=\frac{1}{1!}x^1-\frac{1}{3!}x^3+\frac{1}{5!}x^5-\frac{1}{7!}x^7~\pm...\] so the question would be, how accurate do you need the results?

OpenStudy (amistre64):

the sum of the first 3 terms are accurate to 5 decimal points